# Constacyclic codes with best-known parameters

### v1

@@ 2.3 The case with two kinds of cyclotomic cosets
In this section, we assume that there are exactly two possible values of $l_{i}$, that is, $\{l_{i}:i\in Z_{n,r}\}=\{l,m\}$ for some positive integers $l,m$, where $l<m$ and $l\mid m$.
@@
Write $nr=(q^{m}-1)/s$ for some positive integers $m,s$. Assume $\{l_{i}:i\in Z_{n,r}\}=\{l,m\}$ for some positive integer $l<m$ and $N_{l}>1$. Then,
$\delta_{\lceil N_{l}/2\rceil+1}^{(l)}\geq\delta_{\lfloor N_{l}/2\rfloor+1}^{(l)}>\frac{qN_{m}}{2(q-1)}\cdot r.$
@@
Now, assume $N_{l}$ is even. Then, by Lemma 2.6 we have
@@
Finally, assume that $N_{l}$ is odd and $N_{l}>1$. Considering the case when $q=2$, we have $r=1$, which implies $n\in Z_{n,r}$ and the coset
$C_{n}^{(q,nr)}=\{n\}.$
So, in this case we have $l=1$ and $N_{l}=\gcd(n,1)=1$, which contradicts with $N_{l}>1$. Thus, we must have $q\geq 3$.
@@
By Lemma 2.6, we obtain
$\begin{split}\delta_{\lfloor N_{l}/2\rfloor+1}>&(N_{l}/2-1/2+1-1)\cdot\frac{nr^{2}}{\gcd(nr,q^{l}-1)}\\ =&\left(\frac{\gcd(n,(q^{l}-1)/r)}{2l}-\frac{1}{2}\right)\cdot\frac{nr^{2}}{\gcd(nr,q^{l}-1)}\\ =&\left(\frac{1}{2l}-\frac{r}{2\gcd(nr,q^{l}-1)}\right)nr\\ =&\left(\frac{1}{2l}-\frac{1}{2lN_{l}}\right)nr=\frac{1}{2l}\cdot\frac{N_{l}-1}{N_{l}}\cdot nr.\end{split}$
@@
$\begin{split}\delta_{\lfloor N_{l}/2\rfloor+1}&>\frac{1}{2l}\cdot\frac{N_{l}-1}{N_{l}}\cdot nr\\ &\geq\frac{1}{m}\cdot\frac{q}{2(q-1)}\cdot nr=\frac{q}{q-1}\cdot\frac{nr}{2m}.\end{split}$
@@
Let $q=7$, $n=19$, $r=6$. Then, $m=3$, and the $7$-cyclotomic cosets in $Z_{19,6}$ are:
$C_{1}^{(7,114)}=\{1,7,49\},\quad C_{13}^{(7,114)}=\{13,91,67\},\quad C_{19}^{(7,114)}=\{19\},$
$C_{25}^{(7,114)}=\{25,61,85\},\quad C_{31}^{(7,114)}=\{31,103,37\},$
$C_{43}^{(7,114)}=\{43,73,55\},\quad C_{79}^{(7,114)}=\{79,97,109\}.$
So, $\{l_{i},i\in Z_{19,6}\}=\{1,3\}$, $N_{1}=1$ and $N_{3}=6$. Noticing that
$\delta_{\lfloor N_{1}/2\rfloor+1}^{(1)}=\delta_{1}^{(1)}=19$
and
$\frac{qN_{m}}{2(q-1)}\cdot r=\frac{7\times 6}{12}\times 6=21,$
we have
$\delta_{\lfloor N_{l}/2\rfloor+1}^{(l)}<\frac{qN_{m}}{2(q-1)}\cdot r.$
@@
Write $nr=(q^{m}-1)/s$ for some positive integers $m,s$. Assume that $\{l_{i}:i\in Z_{n,r}\}=\{l,m\}$ for some positive integer $l<m$, $N_{l}=1$ and $2m(q-1)\geq qlr$. Then,
$\delta_{1}^{(l)}>\frac{qN_{m}}{2(q-1)}\cdot r.$
@@
First, assume $q=2$, then we have $r=1$, which implies $n\in Z_{n,r}$ and the coset
$C_{n}^{(q,nr)}=\{n\}.$
So, we have $l=1$ and $N_{l}=\gcd(n,1)=1$. Hence,
@@
Now, assume $q>2$. By Lemma 2.4, we have $N_{l}=\gcd(n,(q^{l}-1)/r)/l$. Then, noticing $N_{l}=1$, we obtain
$\gcd(n,(q^{l}-1)/r)=l.$
Combining this with the condition $2m(q-1)\geq qlr$ and using Lemma 2.6, we have
$\begin{split}\delta_{1}^{(l)}\geq&\frac{nr}{\gcd(nr,q^{l}-1)}\\ =&\frac{n}{l}=\frac{2m}{l}\cdot\frac{n}{2m}\geq\frac{qlr}{l(q-1)}\cdot\frac{n}{2m}=\frac{q}{q-1}\cdot\frac{nr}{2m}.\end{split}$
@@
In Lemma 2.11, when $r=1,2\text{ or }3$, the extra condition
$2m(q-1)\geq qlr$
in fact always holds.
@@
Write $nr=(q^{m}-1)/s$ for some positive integers $m,s$. Assume that $\{l_{i}:i\in Z_{n,r}\}=\{l,m\}$ for some positive integer $l<m$, $N_{l}=1$ and $r\leq 3$. Then,
$\delta_{1}^{(l)}>\frac{qN_{m}}{2(q-1)}\cdot r.$
@@
Combining (2.5) with (2.6), we get
$2m/l\geq\frac{q}{q-1}\cdot r,$
that is,
$2m(q-1)\geq qlr.$
Thus, by Lemma 2.11, we have
$\delta_{1}^{(l)}>\frac{qN_{m}}{2(q-1)}\cdot r.$
@@
Write $nr=(q^{m}-1)/s$ for some positive integers $m,s$. Assume $\{l_{i}:i\in Z_{n,r}\}=\{1,m\}$ with $m\geq 2$. Then, if $N_{1}\geq 2$, we have
@@
$\begin{split}&\lceil N_{1}/2\rceil+m\left(\lceil N_{m}/2\rceil-1\right)\\ &\leq\frac{\gcd(n,(q-1)/r)}{2}+\frac{1}{2}+m\cdot\frac{n-\gcd(n,(q-1)/r)}{2m}-\frac{m}{2}\\ &=\frac{n}{2}-\frac{m-1}{2}.\end{split}$ (2.10)
However, in the set $Z_{n,r}$ the number of elements less than $\frac{nr}{2}+1$ is at least
$\lceil n/2\rceil\geq n/2,$
which contradicts with (2.10) (noticing $m\geq 2$). Hence, (2.9) is true, and thus we get
$\delta_{\lceil N_{1}/2\rceil+1}^{(1)}>\delta_{\lceil N_{m}/2\rceil}^{(m)}.$
@@
$\begin{split}&\delta_{\lfloor N_{1}/2\rfloor+1}^{(1)}\\ \geq&(\lfloor N_{1}/2\rfloor+1-1)\cdot\frac{nr^{2}}{\gcd(nr,q-1)}+\frac{nr}{\gcd(nr,q-1)}\\ =&\left(\frac{\gcd(n,(q-1)/r)}{2}-\frac{1}{2}\right)\cdot\frac{nr^{2}}{\gcd(nr,q-1)}+\frac{nr}{\gcd(nr,q-1)}\\ =&\frac{nr}{2}+\frac{nr(2-r)}{2\gcd(nr,q-1)}\geq\frac{nr}{2},\end{split}$ (2.11)
@@
$\delta_{\lfloor N_{m}/2\rfloor}^{(m)}<\frac{nr}{2}$ (2.12)
@@
$\begin{split}&\lfloor N_{1}/2\rfloor+m\left(\lfloor N_{m}/2\rfloor-1\right)\\ &\leq\frac{\gcd(n,(q-1)/r)}{2}+m\cdot\frac{n-\gcd(n,(q-1)/r)}{2m}-m\\ &=\frac{n}{2}-m.\end{split}$ (2.13)
However, in the set $Z_{n,r}$ the number of elements less than $nr/2$ is at least
$\lfloor n/2\rfloor-1\geq\frac{n}{2}-\frac{3}{2},$
which contradicts with (2.13) (noticing $m\geq 2$). Therefore, (2.12) is true, and so, we obtain
This completes the proof. ∎
@@
Choosing $q=19$, $n=127$ and $r=18$, we have $nr=(q^{m}-1)/s$ with $m=3$ and $s=3$, $\{l_{i},i\in Z_{127,18}\}=\{1,3\}$, $N_{1}=1$ and $N_{3}=42$. Moreover, $\delta_{1}^{(1)}=127$ and $\delta_{21}^{(3)}=451$. So,
$\delta_{\lfloor N_{1}/2\rfloor+1}^{(1)}=\delta_{1}^{(1)}<\delta_{21}^{(3)}.$
@@ 2.4 Cyclotomic values
$\Phi_{m}(x)=\prod_{d\mid m}(x^{d}-1)^{\mu(m/d)},$ (2.14)
where $\mu$ is the Möbius function.
The following result is in [37, Proposition 2.8] about $\Phi_{m}(1)$.
@@
If $m$ has at least two distinct prime divisors, then $\Phi_{m}(1)=1$.
@@
which contradicts with $p\mid\Phi_{m}(q)$. Thus, $\gcd(\Phi_{m}(q),q-1)=1$. This proves the first part of the lemma.
@@
So, we have
$\gcd(\Phi_{m}(q),q-1)=\gcd(q-1,p).$
This proves the second part of the lemma. ∎
@@ 2.5 The BCH bound
The BCH lower bound is an important bound for the minimun distance of BCH codes, which can be extended for constacyclic codes; see [18, Theorem 2.2].
@@ 3 Some general constructions for constacyclic codes
In this section, motivated by [32] and generalizing the constructions in [8, 20, 23, 31, 32], we propose a general construction for $q$-ary constacyclic codes. First we classify the $q$-cyclotomic cosets in $Z_{n,r}$ according to their sizes (that is, two cosets are in the same class if and only if they have the same size), and then we take the defining set $T$ as the union of the first half of cosets in each class.
@@
For simplicity, we still denote $N_{l}=N_{l}^{(q,nr)}$ and $\delta_{i}^{(l)}=\delta_{i}^{(q,nr,l)}$ for any positive integer $l$. By definition, $N_{l}$ is the number of $q$-cyclotomic cosets with $l$ elements in the set $Z_{n,r}$, and $\delta_{i}^{(l)}$ is the $i$-th coset leader (in ascending order) in those $q$-cyclotomic cosets with $l$ elements in $Z_{n,r}$.
Recall that for any $i\in Z_{n,r}$, $l_{i}=|C_{i}^{(q,nr)}|$, which is the size of the $q$-cyclotomic coset generated by $i$ modulo $nr$.
@@
Write $nr=(q^{m}-1)/s$, and assume $\{l_{i}:i\in Z_{n,r}\}=\{l,m\}$ for some positive integer $l<m$. Then, an infinite family of $q$-ary constacyclic $n,n-l\lceil N_{l}/2\rceil-m\lceil N_{m}/2\rceil,\geq\lfloor\frac{qN_{m}}{2(q-1)}\rfloor]$ codes is constructed, where
@@
then an infinite family of constacyclic $n,n-l\lfloor N_{l}/2\rfloor-m\lfloor N_{m}/2\rfloor,\geq\lfloor\frac{qN_{m}}{2(q-1)}\rfloor]$ codes is constructed.
@@
Take the defining set $T$ as the union of the first $\lceil N_{l}/2\rceil$ $q$-cyclotomic cosets contained in $Z_{n,r}$ with $l$ elements and the first $\lceil N_{m}/2\rceil$ $q$-cyclotomic cosets contained in $Z_{n,r}$ with $m$ elements. So, there are $l\lceil N_{l}/2\rceil+m\lceil N_{m}/2\rceil$ integers in $T$, which implies that the dimension of the corresponding constacyclic code $\mathcal{C}$ is
$n-l\lceil N_{l}/2\rceil-m\lceil N_{m}/2\rceil.$
Next, we prove the desired lower bound for the minimun distance $d(\mathcal{C})$. Since $d(\mathcal{C})\geq 2$, we can assume $N_{m}>4$ without loss of generality. When $N_{l}>1$, by Lemma 2.7 and Lemma 2.9, we have
$\delta_{\lceil N_{l}/2\rceil+1}^{(l)}>\frac{qN_{m}}{2(q-1)}\cdot r>1+\left(\frac{qN_{m}}{2(q-1)}-2\right)r$
and
@@
By construction, this means that for any non-negative integer $i\leq\frac{qN_{m}}{2(q-1)}-2$, $1+ir\in T$. So, by Lemma 2.17 we get
$d(\mathcal{C})\geq\left\lfloor\frac{qN_{m}}{2(q-1)}\right\rfloor.$
Similarly, the case when $N_{l}=1$ follows from Lemma 2.9 and Lemma 2.17. So, we prove the first part of the theorem.
Now, take the defining set $T$ as the union of the first $\lfloor N_{l}/2\rfloor$ $q$-cyclotomic cosets contained in $Z_{n,r}$ with $l$ elements and the first $\lfloor N_{m}/2\rfloor$ $q$-cyclotomic cosets contained in $Z_{n,r}$ with $m$ elements. Then, under one of the conditions (1), (2) and (3) and using Lemmas 2.9 and 2.11 and Corollary 2.12, an infinite family of constacyclic $n,n-l\lfloor N_{l}/2\rfloor-m\lfloor N_{m}/2\rfloor,\geq\lfloor\frac{qN_{m}}{2(q-1)}\rfloor]$ codes is constructed. ∎
We remark that the quantity $N_{m}$ in Theorem 3.1 is greater than $cn/\log_{q}n$ for some constant $c$ which does not depend on $m$. This means that the minimum distance $d(\mathcal{C})$ is also greater than $cn/\log_{q}n$.
The following example suggests that when $r=1$, the constructed codes in Theorem 3.1 may be not always BCH codes.
@@
Let $q=5$, $n=26$, and $r=1$. Then, there are eight $5$-cyclotomic cosets modulo $26$:
$C_{1}^{(5,26)}=\{1,5,25,21\},\quad C_{2}^{(5,26)}=\{2,10,24,16\},$
$C_{3}^{(5,26)}=\{3,15,23,11\},\quad C_{4}^{(5,26)}=\{4,20,22,6\},$
$C_{7}^{(5,26)}=\{7,9,19,17\},\quad C_{8}^{(5,26)}=\{8,14,18,12\},$
$C_{13}^{(5,26)}=\{13\},\qquad\qquad C_{26}^{(5,26)}=\{26\}.$
Then, $N_{1}=2,N_{4}=6$. So, in Theorem 3.1, the defining set $T$ is
$C_{1}^{(5,26)}\cup C_{2}^{(5,26)}\cup C_{3}^{(5,26)}\cup C_{13}^{(5,26)},$
and the corresponding code is not a BCH code.
Now, we want to establish a better lower bound for minimum distance in some special cases.
@@
Write $nr=(q^{m}-1)/s$, and assume $\{l_{i}:i\in Z_{n,r}\}=\{1,m\}$ for some positive integer $m>1$. Then, for the first family of $q$-ary constacyclic codes in Theorem 3.1, the lower bound on their minimum distances can be

### v2

@@ 2.3 The case with two kinds of cyclotomic cosets
In this section, we write $nr=(q^{m}-1)/s$ for some positive integers $m,s$. We also assume that there are exactly two possible values of $l_{i}$ (recalling $l_{i}=|C_{i}|$), that is, $\{l_{i}:i\in Z_{n,r}\}=\{l,m\}$ for some positive integers $l,m$, where $l<m$ and $l\mid m$.
@@
Assume that $\{l_{i}:i\in Z_{n,r}\}=\{l,m\}$ for some positive integer $l<m$ and $N_{l}>1$. Then,
$\delta_{\lceil N_{l}/2\rceil+1}^{(l)}\geq\delta_{\lfloor N_{l}/2\rfloor+1}^{(l)}>\frac{qN_{m}}{2(q-1)}\cdot r,$
where $N_{m}=(n-\gcd(n,(q^{l}-1)/r))/m$.
@@
Now, assume that $N_{l}$ is even. Then, by the second lower bound in Lemma 2.6 we have
@@
Finally, assume that $N_{l}$ is odd and $N_{l}>1$. Considering the case when $q=2$, we have $r=1$, which implies $n\in Z_{n,r}$ and the cyclotomic coset $C_{n}=\{n\}$. So, in this case we have $l=1$ and $N_{l}=\gcd(n,1)=1$, which contradicts with $N_{l}>1$. Thus, we must have $q\geq 3$.
@@
By Lemma 2.6 and noticing that $N_{l}$ is odd, we obtain
$\begin{split}\delta_{\lfloor N_{l}/2\rfloor+1}^{(l)}>&(N_{l}/2-1/2+1-1)\cdot\frac{nr^{2}}{\gcd(nr,q^{l}-1)}\\ =&\left(\frac{\gcd(n,(q^{l}-1)/r)}{2l}-\frac{1}{2}\right)\cdot\frac{nr^{2}}{\gcd(nr,q^{l}-1)}\\ =&\left(\frac{1}{2l}-\frac{r}{2\gcd(nr,q^{l}-1)}\right)nr\\ =&\left(\frac{1}{2l}-\frac{1}{2lN_{l}}\right)nr=\frac{1}{2l}\cdot\frac{N_{l}-1}{N_{l}}\cdot nr.\end{split}$
@@
$\begin{split}\delta_{\lfloor N_{l}/2\rfloor+1}^{(l)}>\frac{1}{2l}\cdot\frac{N_{l}-1}{N_{l}}\cdot nr\geq\frac{1}{m}\cdot\frac{q}{2(q-1)}\cdot nr=\frac{q}{q-1}\cdot\frac{nr}{2m}.\end{split}$
@@
Let $q=7$, $n=19$, $r=6$. Then, $m=3$, $s=3$, and the $7$-cyclotomic cosets in $Z_{19,6}$ are: $C_{1}=\{1,7,49\},C_{13}=\{13,67,91\},C_{19}=\{19\},C_{25}=\{25,61,85\},C_{31}=\{31,37,103\},C_{43}=\{43,55,73\},C_{79}=\{79,97,109\}$. So, $\{l_{i},i\in Z_{19,6}\}=\{1,3\}$, $N_{1}=1$ and $N_{3}=6$. Then, $\delta_{\lfloor N_{1}/2\rfloor+1}^{(1)}=\delta_{1}^{(1)}=19$, but $\frac{qN_{m}}{2(q-1)}\cdot r=\frac{7\times 6}{12}\times 6=21>19.$
@@
Assume that $\{l_{i}:i\in Z_{n,r}\}=\{l,m\}$ for some positive integer $l<m$, $N_{l}=1$ and $2m(q-1)\geq qlr$. Then,
$\delta_{1}^{(l)}>\frac{qN_{m}}{2(q-1)}\cdot r,$
where $N_{m}=(n-\gcd(n,(q^{l}-1)/r))/m$.
@@
First, assume $q=2$, then we have $r=1$, which implies $n\in Z_{n,r}$ and the cyclotomic coset $C_{n}=\{n\}$. So, we have $l=1$ and $N_{l}=\gcd(n,1)=1$. Hence,
@@
Now, assume $q>2$. By Lemma 2.4, we have $N_{l}=\gcd(n,(q^{l}-1)/r)/l$. Then, noticing $N_{l}=1$, we obtain $\gcd(n,(q^{l}-1)/r)=l$. Combining this with the condition $2m(q-1)\geq qlr$ and using the first lower bound in Lemma 2.6, we have
$\begin{split}\delta_{1}^{(l)}\geq\frac{nr}{\gcd(nr,q^{l}-1)}=\frac{n}{l}=\frac{2m}{l}\cdot\frac{n}{2m}\geq\frac{qlr}{l(q-1)}\cdot\frac{n}{2m}=\frac{q}{q-1}\cdot\frac{nr}{2m}.\end{split}$
@@
In Lemma 2.11, when $r=1,2\text{ or }3$, the extra condition $2m(q-1)\geq qlr$ in fact always holds.
@@
Assume that $\{l_{i}:i\in Z_{n,r}\}=\{l,m\}$ for some positive integer $l<m$, $N_{l}=1$ and $r\leq 3$. Then,
$\delta_{1}^{(l)}>\frac{qN_{m}}{2(q-1)}\cdot r,$
where $N_{m}=(n-\gcd(n,(q^{l}-1)/r))/m$.
@@
Combining (2.5) with (2.6), we obtain $2m/l\geq\frac{q}{q-1}\cdot r,$ that is, $2m(q-1)\geq qlr.$ Thus, by Lemma 2.11, we get the desired result. ∎
@@
Assume $\{l_{i}:i\in Z_{n,r}\}=\{1,m\}$ with $m\geq 2$. Then, if $N_{1}\geq 2$, we have
@@
$\begin{split}\lceil N_{1}/2\rceil+m\left(\lceil N_{m}/2\rceil-1\right)&\leq\frac{\gcd(n,(q-1)/r)}{2}+\frac{1}{2}+m\cdot\frac{n-\gcd(n,(q-1)/r)}{2m}-\frac{m}{2}\\ &=\frac{n}{2}-\frac{m-1}{2}.\end{split}$ (2.10)
However, in the set $Z_{n,r}$ the number of elements less than $\frac{nr}{2}+1$ is at least $\lceil n/2\rceil\geq n/2,$ which contradicts with (2.10) (noticing $m\geq 2$). Hence, (2.9) is true, and thus we get $\delta_{\lceil N_{1}/2\rceil+1}^{(1)}>\delta_{\lceil N_{m}/2\rceil}^{(m)}.$
@@
$\begin{split}\delta_{\lfloor N_{1}/2\rfloor+1}^{(1)}&\geq(\lfloor N_{1}/2\rfloor+1-1)\cdot\frac{nr^{2}}{\gcd(nr,q-1)}+\frac{nr}{\gcd(nr,q-1)}\\ &=\left(\frac{\gcd(n,(q-1)/r)}{2}-\frac{1}{2}\right)\cdot\frac{nr^{2}}{\gcd(nr,q-1)}+\frac{nr}{\gcd(nr,q-1)}\\ &=\frac{nr}{2}+\frac{nr(2-r)}{2\gcd(nr,q-1)}\geq\frac{nr}{2},\end{split}$ (2.11)
@@
$\delta_{\lfloor N_{m}/2\rfloor}^{(m)}<\frac{nr}{2}.$ (2.12)
@@
$\begin{split}\lfloor N_{1}/2\rfloor+m\left(\lfloor N_{m}/2\rfloor-1\right)&\leq\frac{\gcd(n,(q-1)/r)}{2}+m\cdot\frac{n-\gcd(n,(q-1)/r)}{2m}-m\\ &=\frac{n}{2}-m.\end{split}$ (2.13)
However, in the set $Z_{n,r}$ the number of elements less than $nr/2$ is at least $\lfloor n/2\rfloor-1\geq\frac{n}{2}-\frac{3}{2},$ which contradicts with (2.13) (noticing $m\geq 2$). Therefore, (2.12) is true, and so, we obtain $\delta_{\lfloor N_{1}/2\rfloor+1}^{(1)}>\delta_{\lfloor N_{m}/2\rfloor}^{(m)}.$ This completes the proof. ∎
@@
Choosing $q=19$, $n=127$ and $r=18$, we have $nr=(q^{m}-1)/s$ with $m=3$, $s=3$, $\{l_{i},i\in Z_{127,18}\}=\{1,3\}$, $N_{1}=1$ and $N_{3}=42$. Moreover, $\delta_{1}^{(1)}=127$ and $\delta_{21}^{(3)}=451$. So, $\delta_{\lfloor N_{1}/2\rfloor+1}^{(1)}=\delta_{1}^{(1)}<\delta_{21}^{(3)}=\delta_{\lfloor N_{3}/2\rfloor}^{(m)}.$ In addition, $\delta_{1}^{(1)}$ is also less than the 21-th element 361 in the set $Z_{127,18}$, and so, $\delta_{1}^{(1)}$ is not in the first 21 cyclotomic cosets of size 3 in $Z_{127,18}$.
@@ 2.4 Cyclotomic values
$\Phi_{m}(x)=\prod_{d\mid m}(x^{d}-1)^{\mu(m/d)}.$ (2.14)
The following result is in [44, Proposition 2.8] about $\Phi_{m}(1)$.
@@
If $m$ has at least two distinct prime factors, then $\Phi_{m}(1)=1$.
@@
which contradicts with $p\mid\Phi_{m}(q)$. Thus, $\gcd(\Phi_{m}(q),q-1)=1$.
@@
So, we have $\gcd(\Phi_{m}(q),q-1)=\gcd(q-1,p).$ This proves the second part of the lemma. ∎
@@ 2.5 The BCH bound
The BCH lower bound is an important bound for the minimun distance of BCH codes, which can be extended for constacyclic codes; see [22, Theorem 2.2].
@@
2.6 Equivalence between constacyclic codes and cyclic codes
It is well-known that a $\lambda$-constacyclic code over $\mathbb{F}_{q}$ is monomially equivalent to a cyclic code of the same length over the field $\mathbb{F}_{q}(\alpha)$, where $\alpha$ is an $n$-th root of $\lambda$; see [3, Theorem 2.1]. The following lemma is the special case when $\alpha\in\mathbb{F}_{q}$; see [3, Corollary 2.1 and Lemma 2.1]. Recall that $\lambda\in\mathbb{F}_{q}^{*}$ and $r$ is the multiplicative order of $\lambda$.
Lemma 2.18 ([3]).
If $\gcd(n,q-1)\mid\frac{q-1}{r}$, then a $\lambda$-constacyclic code of length $n$ over $\mathbb{F}_{q}$ is monomially equivalent to a cyclic code of length $n$ over $\mathbb{F}_{q}$.
From Lemma 2.18, we get the following corollary.
Corollary 2.19.
If $\gcd(r,\frac{nr}{\gcd(nr,q-1)})=1$, then a $\lambda$-constacyclic code of length $n$ over $\mathbb{F}_{q}$ is monomially equivalent to a cyclic code of length $n$ over $\mathbb{F}_{q}$.
Proof.
Since $r\mid q-1$, we have $\gcd(r,\frac{nr}{\gcd(nr,q-1)})=\gcd(r,\frac{n}{\gcd(n,(q-1)/r)})$. So, by assumption we have $\gcd(r,\frac{n}{\gcd(n,(q-1)/r)})=1$. Then, we get
$\gcd(n,q-1)=\gcd(n,r\cdot(q-1)/r)$ $=\gcd(n,\frac{q-1}{r})\cdot\gcd(\frac{n}{\gcd(n,\frac{q-1}{r})},r)$
$=\gcd(n,\frac{q-1}{r}),$
which implies $\gcd(n,q-1)\mid\frac{q-1}{r}$. Hence, the desired result follows from Lemma 2.18. ∎
We remark that the condition $\gcd(r,\frac{nr}{\gcd(nr,q-1)})=1$ in Corollary 2.19 is equivalent to that $N_{1}>0$ (that is, there exists a $q$-cyclotomic coset with one element in $Z_{n,r}$ modulo nrnr); see Lemma 2.4.
@@ 3 Some general constructions for constacyclic codes
In this section, motivated by [39] and generalizing the constructions in [9, 25, 29, 38, 39], to construct a $q$-ary constacyclic code, we first classify the $q$-cyclotomic cosets in $Z_{n,r}$ according to their sizes, and then we take the defining set $T$ as the union of the first half of cyclotomic cosets in each class (sorted in ascending order by their coset leaders).
@@
For simplicity, we still denote $N_{l}=N_{l}^{(q,n,r)}$ and $\delta_{i}^{(l)}=\delta_{i}^{(q,n,r,l)}$. By definition, $N_{l}$ is the number of $q$-cyclotomic cosets with $l$ elements in the set $Z_{n,r}$, and $\delta_{i}^{(l)}$ is the $i$-th coset leader (in ascending order) in those $q$-cyclotomic cosets with $l$ elements in $Z_{n,r}$.
Recall that for any $i\in Z_{n,r}$, $l_{i}=|C_{i}|$, which is the size of the $q$-cyclotomic coset generated by $i$ modulo $nr$ (we denote $C_{i}=C_{i}^{(q,nr)}$).
@@
Write $nr=(q^{m}-1)/s$, and assume $\{l_{i}:i\in Z_{n,r}\}=\{l,m\}$ for some positive integer $l<m$. Then, an infinite family of $q$-ary $n,n-l\lceil N_{l}/2\rceil-m\lceil N_{m}/2\rceil,\geq\lfloor\frac{qN_{m}}{2(q-1)}\rfloor]$ constacyclic codes is constructed, where
@@
then an infinite family of $q$-ary $n,n-l\lfloor N_{l}/2\rfloor-m\lfloor N_{m}/2\rfloor,\geq\lfloor\frac{qN_{m}}{2(q-1)}\rfloor]$ constacyclic codes is constructed.
@@
First, take the defining set $T$ as the union of the first $\lceil N_{l}/2\rceil$ $q$-cyclotomic cosets contained in $Z_{n,r}$ with $l$ elements and the first $\lceil N_{m}/2\rceil$ $q$-cyclotomic cosets contained in $Z_{n,r}$ with $m$ elements. So, there are $l\lceil N_{l}/2\rceil+m\lceil N_{m}/2\rceil$ integers in $T$, which implies that the dimension of the corresponding constacyclic code $\mathcal{C}$ is $n-l\lceil N_{l}/2\rceil-m\lceil N_{m}/2\rceil.$
Next, we prove the desired lower bound for the minimun distance $d(\mathcal{C})$. Since $d(\mathcal{C})\geq 2$, we can assume the desired lower bound $\lfloor\frac{qN_{m}}{2(q-1)}\rfloor\geq 2$ without loss of generality. When $N_{l}>1$, by Lemmas 2.7 and 2.9, we have
$\delta_{\lceil N_{l}/2\rceil+1}^{(l)}>\frac{qN_{m}}{2(q-1)}\cdot r>1+\left(\frac{qN_{m}}{2(q-1)}-2\right)r,$
@@
By construction, this means that for any non-negative integer $i\leq\frac{qN_{m}}{2(q-1)}-2$, we have $1+ir\in T$. So, by Lemma 2.17 we get $d(\mathcal{C})\geq\left\lfloor\frac{qN_{m}}{2(q-1)}\right\rfloor.$ Similarly, the case when $N_{l}=1$ follows from Lemmas 2.7 and 2.17. So, we prove the first part of the theorem.
Now, take the defining set $T$ as the union of the first $\lfloor N_{l}/2\rfloor$ $q$-cyclotomic cosets contained in $Z_{n,r}$ with $l$ elements and the first $\lfloor N_{m}/2\rfloor$ $q$-cyclotomic cosets contained in $Z_{n,r}$ with $m$ elements. Then, under one of the conditions (1), (2) and (3) and using Lemmas 2.7, 2.9 and 2.11 and Corollary 2.12, an infinite family of constacyclic $n,n-l\lfloor N_{l}/2\rfloor-m\lfloor N_{m}/2\rfloor,\geq\lfloor\frac{qN_{m}}{2(q-1)}\rfloor]$ codes is constructed. ∎
Remark 3.2.
In Theorem 3.1, if moreover $l=1$, then $N_{1}>0$; and so, by Lemma 2.4 we have $\gcd(r,\frac{nr}{\gcd(nr,q-1)})=1$. Then, by Corollary 2.19 we know that the constacyclic codes in Theorem 3.1 are monomially equivalent to some cyclic codes over $\mathbb{F}_{q}$.
We also remark that Theorem 3.1 includes the codes constructed in [9, Theorems 10, 15 and 16] and also improves the lower bounds on minimum distances there.
We notice that in Theorem 3.1, the quantity $N_{m}$ satisfies $N_{m}\geq\frac{n}{2m}=\frac{n}{2\log_{q}(nrs+1)}.$ This means that when $s\leq n^{10}$, the minimum distance $d(\mathcal{C})$ is greater than $cn/\log_{q}n$ for some positive constant $c$ depending only on $q$.
We emphasize that the codes in Theorem 3.1 are not always BCH constacyclic codes; see the following example for the case when $r=1$.
@@
Let $q=5$, $n=26$, and $r=1$. Then, the $5$-cyclotomic cosets modulo $26$ are: $C_{1}=\{1,5,21,25\},C_{2}=\{2,10,16,24\},C_{3}=\{3,11,15,23\},C_{4}=\{4,6,20,22\},C_{7}=\{7,9,17,19\},C_{8}=\{8,12,14,18\},C_{13}=\{13\},C_{26}=\{26\}.$ Then, $N_{1}=2,N_{4}=6$. So, in Theorem 3.1, the defining set is $C_{1}\cup C_{2}\cup C_{3}\cup C_{13}$, and the relevant code is not a BCH code.
Next, by extending the strategy in [39, Theorem 10], we establish a better lower bound for the minimum distance in some special cases.
@@
Write $nr=(q^{m}-1)/s$, and assume $\{l_{i}:i\in Z_{n,r}\}=\{l,m\}$ for some positive integer $l<m$. Then, if either $l=1$ or $N_{l}=1$, for the first family of $q$-ary constacyclic codes in Theorem 3.1, the lower bound on their minimum distances can be